A mixture of CaCl2 and NaCl weighing 4.44g is treated with sodium carbonate solution of precipitate all the Ca2+ ions as calcium carbonate. The calcium carbonate so obtained is heated strongly to get 0.56g of CaO. The percentage of NaCl in the mixture is [ans:75]
 

Grade:11

1 Answers

Arun
5 years ago
Weight of CaO = 0.56 g
0.56 g of CaO contains 0.40 g of calcium. Mol. weight of CaCl2 = 40 + 71 = 111
40 g of calcium present in 111 g of CaCl2
∴ 0.4 g of calcium present in 0.40 x 111/40 = 1.11 g of CaCl2
Amount of NaCl present in mixture = 4.44 – 1.11 = 3.33 g
∴ Percentage of NaCl = 3.33 /4.44x 100 = 75